OpenStudy (anonymous):

The formula for degrees Celsius is: C = 5/9 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 4 degrees Celsius is. <3 : )

OpenStudy (anonymous):

C = 5/9 (F − 32) C/(5/9) = F-32 9C/5 = F-32 F = 9C/5 + 32

OpenStudy (anonymous):

F = 36/5 + 32 F = 39.2